Advertisements
Advertisements
प्रश्न
Three integers are chosen at random from the first 20 integers. The probability that their product is even is
विकल्प
2/19
3/29
17/19
4/19
Advertisements
उत्तर
17/19
Number of ways in which we can choose three distinct integers from 20 integers = 20C3= 1140
We know that, if we take three odd numbers, there product will always be an odd number.
Out of 20 consecutive integers, 10 are even and 10 are odd integers.
Number of ways in which we can choose three distinct odd integers from 10 odd integers= 10C3 = 120
P(product is even) = 1 - P(product is odd)
= \[1 - \frac{120}{1140} = \frac{1140 - 120}{1140} = \frac{1020}{1140} = \frac{17}{19}\]
APPEARS IN
संबंधित प्रश्न
A letter is chosen at random from the word ‘ASSASSINATION’. Find the probability that letter is
- a vowel
- an consonant
A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P(not A).
In an entrance test that is graded on the basis of two examinations, the probability of a randomly chosen student passing the first examination is 0.8 and the probability of passing the second examination is 0.7. The probability of passing at least one of them is 0.95. What is the probability of passing both?
In a certain lottery, 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy one ticket.
Out of 100 students, two sections of 40 and 60 are formed. If you and your friend are among the 100 students, what is the probability that
- you both enter the same sections?
- you both enter the different sections?
A dice is thrown. Find the probability of getting:
2 or 4
A dice is thrown. Find the probability of getting a multiple of 2 or 3.
In a simultaneous throw of a pair of dice, find the probability of getting:
8 as the sum
In a simultaneous throw of a pair of dice, find the probability of getting a doublet
In a simultaneous throw of a pair of dice, find the probability of getting a doublet of prime numbers
In a simultaneous throw of a pair of dice, find the probability of getting a doublet of odd numbers
In a simultaneous throw of a pair of dice, find the probability of getting a sum greater than 9
In a simultaneous throw of a pair of dice, find the probability of getting an even number on first
In a simultaneous throw of a pair of dice, find the probability of getting odd number on the first and 6 on the second
Three coins are tossed together. Find the probability of getting exactly two heads
Three coins are tossed together. Find the probability of getting at least two heads
Three coins are tossed together. Find the probability of getting at least one head and one tail.
Two dice are thrown. Find the odds in favour of getting the sum 5.
What are the odds in favour of getting a spade if the card drawn from a well-shuffled deck of cards? What are the odds in favour of getting a king?
A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is white .
A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is white and odd numbered .
A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is even numbered
A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is red or even numbered.
Fill in the blank in the table:
| P (A) | P (B) | P (A ∩ B) | P(A∪ B) |
| 0.35 | .... | 0.25 | 0.6 |
There are three events A, B, C one of which must and only one can happen, the odds are 8 to 3 against A, 5 to 2 against B, find the odds against C
One of the two events must happen. Given that the chance of one is two-third of the other, find the odds in favour of the other.
The probability that a leap year will have 53 Fridays or 53 Saturdays is
A person write 4 letters and addresses 4 envelopes. If the letters are placed in the envelopes at random, then the probability that all letters are not placed in the right envelopes, is
Two dice are thrown simultaneously. The probability of getting a pair of aces is
A box contains 10 good articles and 6 with defects. One item is drawn at random. The probability that it is either good or has a defect is
One mapping is selected at random from all mappings of the set A = {1, 2, 3, ..., n} into itself. The probability that the mapping selected is one to one is
In a certain lottery 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy two tickets.
